# Newton’s method for nonlinear mappings into vector bundles

### v1

@@ 4. Newton-Differentiability and Strict Differentiability
In linear spaces it is well known, and also not hard to verify, that continuous differentiability implies Newton-differentiability, as well as strict differentiability. Since the chart representation of a $C^{1}$-mapping $f:\mathcal{X}\to\mathcal{Y}$ is a $C^{1}$-mapping between vector spaces, we can conclude, that $f$ is also Newton-differentiable and strictly differentiable.
In linear spaces we have the following chain rule for the Newton-differentiability [10] and strict differentiability, which we will use to verify that our definition is independent of the used charts:
Lemma 4.3.
Let $\mathbb{X},\;\mathbb{Y}$ and $\mathbb{Z}$ be normed linear spaces. Consider $f:\mathbb{X}\to\mathbb{Y}$ and $g:\mathbb{Y}\to\mathbb{Z}$.
  1. (i)
    Let $f$ and $g$ be Newton-differentiable with respect to Newton-derivatives $f^{\prime}$ and $g^{\prime}$ in open sets $V$ and $U$, respectively, with $U\subset\mathbb{X}$, $f(U)\subset V\subset\mathbb{Y}$. Assume that $f$ is locally Lipschitz continuous and $g^{\prime}$ is locally bounded. Then $g\circ f:\mathbb{X}\to\mathbb{Z}$ is Newton-differentiable with respect to a Newton-derivative $g^{\prime}(f(\cdot))f^{\prime}(\cdot):\mathbb{X}\to\mathbb{Z}$.
  2. (ii)
    Let $f$ and $g$ be strictly differentiable with strict derivative $f^{\prime}$ and $g^{\prime}$ in open sets $V$ and $U$, respectively, with $U\subset\mathbb{X}$, $f(U)\subset V\subset\mathbb{Y}$. Then $g\circ f:\mathbb{X}\to\mathbb{Z}$ is strictly differentiable with strict derivative $g^{\prime}(f(\cdot))f^{\prime}(\cdot):\mathbb{X}\to\mathbb{Z}$.
Proof.
Let $x,y,x_{0}\in U$. Consider the remainder term
$\|g^{\prime}(f(x_{0}))f^{\prime}(x_{0})(x-y)-(g\circ f(x)-g\circ f(y))\|_{\mathbb{Z}}$ (6)
$=\|g^{\prime}(f(x_{0}))(f(x)-f(y))-(g(f(x))-g(f(y)))+R(x,y,x_{0})\|_{\mathbb{Z}}$
where $R(x,y,x_{0}):=g^{\prime}(f(x_{0}))f^{\prime}(x_{0})(x-y)-g^{\prime}(f(x_{0}))(f(x)-f(y))$.
Since $g^{\prime}$ is locally bounded in the Newton-differentiable case and $x_{0}$ is fixed in the strict differentiable case, there exists a local constant $C>0$ such that
$\|g^{\prime}(f(x_{0}))\|_{\mathbb{Y}\to\mathbb{Z}}\leq C.$
Thus, we obtain
$\|R(x,y,x_{0})\|_{\mathbb{Z}}$ $=\|g^{\prime}(f(x_{0}))\left[f^{\prime}(x_{0})(x-y)-(f(x)-f(y)\right]\|_{\mathbb{Z}}$
$\leq C\cdot\|f^{\prime}(x_{0})(x-y)-(f(x)-f(y))\|_{\mathbb{Y}}$
Hence, we can estimate (6) by
$\|g^{\prime}(f(x_{0}))f^{\prime}(x_{0})(x-y)-(g\circ f(x)-g\circ f(y))\|_{\mathbb{Z}}$ (7)
$\leq\|g^{\prime}(f(x_{0}))(f(x)-f(y))-(g(f(x))-g(f(y)))\|_{\mathbb{Z}}+C\cdot\|f^{\prime}(x_{0})(x-y)-(f(x)-f(y))\|_{\mathbb{Y}}.$
Now consider the different cases:
(i): Set $x=x_{0}$. Let $y$ be fixed. Newton-differentiability of $f$ at $x$ yields
$\|f^{\prime}(x)(x-y)-(f(x)-f(y))\|_{\mathbb{Y}}=o(\|x-y\|_{\mathbb{X}})$
as $x\to y$. Since $g$ is Newton-differentiable at $f(x)\in V$, we get that
$\|g^{\prime}(f(x))(f(x)-f(y))-(g(f(x))-g(f(y)))\|_{\mathbb{Z}}=o(\|f(x)-f(y)\|_{\mathbb{Y}}).$
By local Lipschitz-continuity of $f$, we obtain that $\|f(x)-f(y)\|_{\mathbb{Y}}=\mathcal{O}(\|x-y\|_{\mathbb{X}})$ as $x\to y$. Using this in (7), we get for the remainder term
$\|g^{\prime}(f(x))f^{\prime}(x)(x-y)-(g\circ f(x)-g\circ f(y))\|_{\mathbb{Z}}=o(\|x-y\|_{\mathbb{X}})$
as $x\to y$, and thus Newton-differentiability of $g\circ f$ with respect to $g^{\prime}(f(\cdot))f^{\prime}(\cdot)$.
(ii): Let $\varepsilon>0$ and $x_{0}\in X$ be fixed. Using the strict differentiability of $f$ at $x_{0}$, there exists a neighborhood $W\subset U$ such that for $x,y\in W$:
$\|f^{\prime}(x_{0})(x-y)-(f(x)-f(y))\|_{\mathbb{Y}}<\frac{\varepsilon}{2C}\|x-y\|_{\mathbb{X}}.$
Thus, (7) yields
$\|g^{\prime}(f(x_{0}))f^{\prime}(x_{0})(x-y)-(g\circ f(x)-g\circ f(y))\|_{\mathbb{Z}}$
$\leq\|g^{\prime}(f(x_{0}))(f(x)-f(y))-(g(f(x))-g(f(y)))\|_{\mathbb{Z}}+\frac{\varepsilon}{2}\|x-y\|_{\mathbb{X}}.$
Using that $\|f^{\prime}(x_{0})\|_{\mathbb{X}\to\mathbb{Y}}$ is finite and $f$ is strictly differentiable at $x_{0}$, for every $\tilde{\varepsilon}>0$ there is a neighborhood such that
$\|f(x)-f(y)\|_{\mathbb{Y}}$ $\leq\|f(x)-f(y)-f^{\prime}(x_{0})(x-y)\|_{\mathbb{Y}}+\|f^{\prime}(x_{0})(x-y)\|_{\mathbb{Y}}$ (8)
$<\tilde{\varepsilon}\|x-y\|_{\mathbb{X}}+\|f^{\prime}(x_{0})\|_{\mathbb{X}\to\mathbb{Y}}\|x-y\|_{\mathbb{X}}$
$\leq\widetilde{C}\|x-y\|_{\mathbb{X}}.$
Since $g$ is strictly differentiable at $f(x_{0})$, we obtain by possibly shrinking $W$ that
$\|g^{\prime}(f(x_{0}))(f(x)-f(y))-(g(f(x))-g(f(y)))\|_{\mathbb{Z}}<\frac{\varepsilon}{2\widetilde{C}}\|f(x)-f(y)\|_{\mathbb{Y}}.$
Using (8), we can in summary estimate the remainder term by
$\|g^{\prime}(f(x_{0}))f^{\prime}(x)(x-y)-(g\circ f(x)-g\circ f(y))\|_{\mathbb{Z}}<\varepsilon\|x-y\|_{\mathbb{X}}$
and thus $g\circ f$ is strictly differentiable with strict derivative $g^{\prime}(f(\cdot))f^{\prime}(\cdot)$. ∎
Proposition 4.4.
Newton-differentiability and strict differentiability of $f:\mathcal{X}\to\mathcal{Y}$ are independent of the choice of charts of $\mathcal{X}$ and $\mathcal{Y}$.
Proof.
$\psi_{j}\circ f\circ\phi_{j}^{-1}=T_{ij}^{\mathcal{X}}\circ(\psi_{i}\circ f\circ\phi_{i}^{-1})\circ T_{ji}^{\mathcal{Y}}.$
$T_{ji}^{\mathcal{Y}}$ and $T_{ij}^{\mathcal{X}}$ are continuously differentiable and thus Newton-differentiable, as well as strictly differentiable. By applying the chain rule twice, we obtain Newton- or strict differentiability of $\psi_{j}\circ f\circ\phi_{j}^{-1}$. ∎
@@ 5. Local convergence of Newton’s method
Similar to the case of linear spaces we will now study the local convergence of Newton’s method. This will require some preparation. To formulate local superlinear convergence requires a metric on $\mathcal{X}$ to measure the distance of iterates. In our general setting we are considering Banach manifolds which may not be endowed with a Riemannian metric. Thus, we will discuss a slightly more general way to define a metric on a manifold. Second, we will derive a geometric criterion on Newton- and strict differentiability. This will be especially useful later to connect our a-priori analysis of local convergence with some computable a-posteriori quantities, which are defined in the spirit of Deuflhard [9].
@@ 5.1. Local norms and a metric on manifolds
In Riemannian geometry, manifolds are equipped with a Riemannian metric, based on inner products on the tangent bundle [13, VII §6]. However, in the case of linear spaces, it is known that the convergence theory of Newton’s method is not restricted to Hilbert spaces, but can be applied in general Banach spaces. Hence, we will present a slightly more general notion of a metric, replacing the usual inner products by general norms. In the infinite dimensional case, some care has to be taken, concerning continuity properties.
@@
Consider a function $f:\mathcal{E}\to\mathbb{R}$. We call $f$ fibrewise uniformly continuous around $D\subset E_{y}$, if for any local trivialization $\tau$, the function
@@
For all $\varepsilon>0$ there exists a neighborhood $V\subset\mathcal{M}$ of $y$ and $\delta>0$, such that for all $(y,e_{\tau})\in\tau(D)$ we have the estimate:
$|f_{\tau}(y,e_{\tau})-f_{\tau}(\xi,\eta_{\tau})|<\varepsilon\quad\forall(\xi,\eta_{\tau})\in p^{-1}(U):\xi\in V,\|\eta_{\tau}-e_{\tau}\|_{\mathbb{E}}<\delta.$
It is not hard to show that this notion is independent of the choice of trivialization, taking into account that transition mappings are linear isomorphisms on $\mathcal{E}$.
Clearly, fibrewise uniform continuity is a stronger condition than continuity, since $\delta$ can only depend on $y$ and not on $e_{\tau}$. In particular, we conclude, setting $\eta_{\tau}=e_{\tau}$:
$\sup_{e_{\tau}\in\tau(D)}|f_{\tau}(y,e_{\tau})-f_{\tau}(\xi,e_{\tau})|<\varepsilon\quad\forall\xi\in V.$ (9)
On a vector bundle $p:\mathcal{E}\to\mathcal{M}$ we can define a continuous norm as follows:
@@
Consider a function $\|\cdot\|:\mathcal{E}\to[0,\infty[$ with the following properties:
@@
    For each $y\in\mathcal{M}$ the restriction $\|\cdot\|_{y}$ to $E_{y}$ is a norm on $E_{y}$ which is equivalent to the $\|\cdot\|_{\mathbb{E}}$-norm that makes $\mathbb{E}$ a Banach space. Thus, $(E_{y},\|\cdot\|_{y})$ is a Banach space. We denote by $\mathbb{S}_{y}\subset E_{y}$ the unit sphere with respect to $\|\cdot\|_{y}$.
@@
Let $y\in\mathcal{M}$ and $\|\cdot\|:\mathcal{E}\to[0,\infty)$ be a continuous norm. Consider its representation $\|\cdot\|_{\tau}$ in a local trivialization around $y\in\mathcal{M}$. Then for every $01>\varepsilon>$ 0 there is a neighborhood $V$ of $y$, such that
$(1-\varepsilon)\|e_{\tau}\|_{\tau,y}\leq\|e_{\tau}\|_{\tau,\xi}\leq(1+\varepsilon)\|e_{\tau}\|_{\tau,y}\quad\forall\xi\in V,e_{\tau}\in\mathbb{E}.$
@@
To define a metric on a manifold $\mathcal{X}$ we will consider a continuous norm on the tangent bundle $T\mathcal{X}$. The following integral over a (piecewise) $C^{1}$-curve $\alpha:[a,b]\rightarrow\mathcal{X}$ is well defined, because the integrand is (piecewise) continuous on $[a,b]$:
@@
The value of this integral can be interpreted as the length of the curve $\alpha$. The following derivations are analogous to the construction of a Riemannian distance (cf. e.g. [13, VII §6]), and generalize them from the case of a Hilbert space norm to the case of a Banach space norm.
@@
In the following we will always assume that $\mathcal{X}$ is connected. Then the axioms of a metric follow readily. Symmetry $d(x,y)=d(y,x)$ and non-negativity, $d(x,y)\geq 0$, are obvious from the definition. It is known from topology that on a connected manifold, two points can always be connected by a piecewise smooth curve [12], and thus $d(x,y)<\infty\ \forall x,y\in\mathcal{X}$. The triangle inequality follows straightforwardly from the fact that the length of curves adds up if they are concatenated. It thus remains to show positive definiteness of $d$, which is the only result that requires fibrewise uniform continuity.
To this end, we first prove some results which relates the mapping $d$ to the local norms. Let $(U,\phi)$ be a local chart at $x_{\star}\in\mathcal{X}$ and $\Phi:\pi^{-1}(U)\to U\times\mathbb{X}$ be the corresponding natural vector bundle chart on the tangent bundle $\pi:T\mathcal{X}\to\mathcal{X}$. We define the open ball with radius $r>0$ around 0 with respect to the local norm $\|\cdot\|_{x_{\star}}$ (with representation $\|\cdot\|_{\Phi,x_{\star}}$) as $rB_{x_{\star}}:=\{v\in\mathbb{X}:\|v\|_{\Phi,x_{\star}}<r\}$.
@@
Let $x_{\star}\in\mathcal{X}$ and $(U,\phi)$ a local chart, such that w.l.o.g. $\phi(x_{\star})=0$. Let $x,y\in U$ with representatives $x_{\phi}$ and $y_{\phi}$ in the chart. Let $01>\varepsilon>$ 0. Then there exists a radius $r=r(\varepsilon)>0$ such that for every $C^{1}$-curve $\alpha:[a,b]\to\mathcal{X}$ connecting $x$ and $y$ the following holds:
@@
    If the image $\alpha_{\phi}:[a,b]\to\mathbb{X}$ lies in the open ball $rB_{x_{\star}}$, in particular $x_{\phi},\;y_{\phi}\in rB_{x_{\star}}$, then the following properties hold:
@@
    If $\alpha_{\phi}$ is leaving the ball $rB_{x_{\star}}$, i.e. $\alpha_{\phi}(a)=x_{\phi}\in rB_{x_{\star}}$ but $\alpha_{\phi}\not\subset rB_{x_{\star}}$, then it holds that
    $L(\alpha)\geq(1-\varepsilon)(r-\|x_{\phi}\|_{\Phi,x_{\star}}).$
@@
$(1-\varepsilon)\|v\|_{\Phi,x_{\star}}\leq\|v\|_{\Phi,\xi}\leq(1+\varepsilon)\|v\|_{\Phi,x_{\star}}\;\forall\xi\in V,v\in\mathbb{X}.$
The image of $V$ in the chart contains a $\|\cdot\|_{\Phi,x_{\star}}$-ball $rB_{x_{\star}}$ around $\phi(x_{\star})=0$ with radius $r=r(\varepsilon)>0$. Let $x,y\in U$ and $\alpha:[a,b]\to\mathcal{X}$ be a $C^{1}$-curve connecting $x$ and $y$.
@@
    Assume that the image $\alpha_{\phi}$ lies in $rB_{x_{\star}}$. Using that the curve integral in vector spaces is always greater than or equal to the distance of the end points, we can estimate the length of $\alpha$ by
@@
    Choosing the connecting line of $x_{\phi}$ and $y_{\phi}$ for $\alpha_{\phi}$ this yields $d(x,y)\leq(1+\varepsilon)\|x_{\phi}-y_{\phi}\|_{\Phi,x_{\star}}$.
@@
    If $\alpha_{\phi}$ leaves the open ball $rB_{x_{\star}}$, there must be a first intersection point $s_{\phi}=\alpha_{\phi}(c)$, $c\in[a,b]$, with the sphere $r\mathbb{S}_{x_{\star}}\subset\mathbb{X}$ around 00. Let $\widetilde{\alpha}_{\phi}$ be the part of $\alpha_{\phi}$ connecting $x_{\phi}$ and $s_{\phi}$. Then by the norm equivalence and the inverse triangle inequality:
@@
    $\geq(1-\varepsilon)(\|s_{\phi}\|_{\Phi,x_{\star}}-\|x_{\phi}\|_{\Phi,x_{\star}})=(1-\varepsilon)(r-\|x_{\phi}\|_{\Phi,x_{\star}}).$
    Thus, we get $L(\alpha)\geq(1-\varepsilon)(r-\|x_{\phi}\|_{\Phi,x_{\star}})$.
@@
Let $01>\varepsilon>$ 0. Then there is a neighborhood $W\subseteq U$ of $x_{\star}$ such that
$(1-\varepsilon)\cdot\|x_{\phi}-y_{\phi}\|_{\Phi,x_{\star}}\leq d(x,y)\leq(1+\varepsilon)\cdot\|x_{\phi}-y_{\phi}\|_{\Phi,x_{\star}}\ \forall x,y\in W,$
@@
W.l.o.g. assume $\phi(x_{\star})=0$. By Lemma 5.4, we directly obtain
$d(x,y)\leq(1+\varepsilon)\|x_{\phi}-y_{\phi}\|_{\Phi,x_{\star}}\;\forall x,y\in\phi^{-1}(rB_{x_{\star}}),$
@@
$\alpha_{\phi}\subset rB_{x_{\star}}$ $\Rightarrow L(\alpha)\geq(1-\varepsilon)\|x_{\phi}-y_{\phi}\|_{\Phi,x_{\star}},$
$\alpha_{\phi}\not\subset rB_{x_{\star}}$ $\Rightarrow L(\alpha)\geq(1-\varepsilon)(r-\|x_{\phi}\|_{\Phi,x_{\star}}).$
@@
$d(x,y)\geq(1-\varepsilon)\cdot\min\{\|x_{\phi}-y_{\phi}\|_{\Phi,x_{\star}},r-\|x_{\phi}\|_{\Phi,x_{\star}}\}.$ (10)
Choosing $W:=\phi^{-1}(\frac{r}{3}B_{x_{\star}})$ yields $d(x,y)\geq(1-\varepsilon)\|x_{\phi}-y_{\phi}\|_{\Phi,x_{\star}}$ and thus our claim. ∎
@@
Let $\mathcal{X}$ be a connected $C^{1}$ Banach manifold with the Hausdorff property. Then $d$ defines a metric on $\mathcal{X}$. Moreover, the induced topology coincides with the topology on $\mathcal{X}$ induced by the atlas.
@@
Since $\mathcal{X}$ satisfies the Hausdorff property and $x\neq y$ holds, there is a $\|\cdot\|_{\Phi,x}$-ball $rB_{x}$ around 00 with radius $r>0$ such that $y\not\in\phi^{-1}(rB_{x})$. Possibly decreasing $r$ first, we obtain by (10)
@@
Moreover, for $y\in\phi^{-1}(r\mathbb{S}_{x})$ we obtain by Prop. 5.5
@@
Now consider a retraction $R:T\mathcal{X}\rightarrow\mathcal{X}$. For $C^{1}$-retractions of this kind it is well known that the following property holds [1] by the inverse function theorem.

### v2

@@ 4. Newton-Differentiability and Strict Differentiability
Since the chart representation of a $C^{1}$-mapping $F:\mathcal{X}\to\mathcal{Y}$ is a $C^{1}$-mapping between linear spaces, we can conclude, that such $F$ is also Newton-differentiable and strictly differentiable.
@@ 5. Local convergence of Newton’s method
In this section we will study local convergence of Newton’s method for mappings into vector bundles. This will require some preparation. Formulating local superlinear convergence requires a metric on $\mathcal{X}$ to measure the distance of iterates. In our general setting we are considering Banach manifolds which may not be endowed with a Riemannian metric. Thus, we will discuss a slightly more general way to define a metric on a manifold. Second, we will derive a geometric criterion on Newton- and strict differentiability. This will be especially useful later to connect our a-priori analysis of local convergence with some computable a-posteriori quantities, which are defined in the spirit of Deuflhard [9]. With these tools we can then give criteria for local superlinear convergence of Newton’s method.
@@ 5.1. Local norms and a metric on manifolds
In Riemannian geometry, manifolds are equipped with a Riemannian metric, based on inner products on the tangent bundle [14, VII §6]. However, in the case of linear spaces, it is known that the convergence theory of Newton’s method is not restricted to Hilbert spaces, but can be applied in general Banach spaces. Hence, we will present a slightly more general notion of a metric, replacing the usual inner products by general norms. In the infinite dimensional case, some care has to be taken, concerning continuity properties.
@@
Consider a function $f:\mathcal{E}\to\mathbb{R}$ on a vector bundle $p:\mathcal{E}\to\mathcal{Y}$. We call $f$ fibrewise uniformly continuous around $D\subset E_{y}$, $y\in\mathcal{Y}$, if for any local trivialization $\tau$, the function
@@
For all $\varepsilon>0$ there exists a neighborhood $U\subset\mathcal{Y}$ of $y$ and $\delta>0$, such that for all $(y,e_{\tau})\in\tau(D)$ we have the estimate:
$|f_{\tau}(y,e_{\tau})-f_{\tau}(\hat{y},\hat{e}_{\tau})|<\varepsilon\quad\forall(\hat{y},\hat{e}_{\tau})\in p^{-1}(U):\hat{y}\in U,\|\hat{e}_{\tau}-e_{\tau}\|_{\mathbb{E}}<\delta.$
It is not hard to show that this notion is independent of the choice of trivialization, taking into account that transition mappings are fibrewise linear isomorphisms.
Clearly, fibrewise uniform continuity is a stronger condition than continuity, since $U$ and $\delta$ only depend on $y$ and not on $e_{\tau}$. In particular, we conclude, setting $\hat{e}_{\tau}=e_{\tau}$:
$\sup_{e_{\tau}\in\tau(D)}|f_{\tau}(y,e_{\tau})-f_{\tau}(\hat{y},e_{\tau})|<\varepsilon\quad\forall\hat{y}\in U.$ (10)
On a vector bundle $p:\mathcal{E}\to\mathcal{Y}$ we can define a continuous norm as follows:
@@
Consider a function $\|\cdot\|:\mathcal{E}\to[0,\infty)$ with the following properties:
@@
    For each $y\in\mathcal{Y}$ the restriction $\|\cdot\|_{y}$ to $E_{y}$ is a norm on $E_{y}$ which is equivalent to the $\|\cdot\|_{\mathbb{E}}$-norm that makes $\mathbb{E}$ a Banach space. Thus, $(E_{y},\|\cdot\|_{y})$ is a Banach space. We denote by $\mathbb{S}_{y}\subset E_{y}$ the unit sphere with respect to $\|\cdot\|_{y}$.
@@
Example 5.3.
Let $\mathbb{E}=L_{q}(\Omega)$ for $1\leq q\leq\infty$ and some measurabe set $\Omega$. Further let $s:\mathcal{Y}\times\Omega\to(0,\infty)$ be a function, such that each $s(y,\cdot):\Omega\to(0,\infty)$ is measurable and bounded from above and below by positive constants (which may depend on $y$). Then for each $y\in\mathcal{Y}$, $\|v\|_{y}:=\|s(y,\cdot)v\|_{L_{q}}$ defines a norm that is equivalent to $\|\cdot\|_{L_{q}}$. If we have in addition that $\lim_{\hat{y}\to y}\|s(y,\cdot)-s(\hat{y},\cdot)\|_{\infty}=0$ for all $y\in\mathcal{Y}$, then this norm is continuous in the sense of Definition 5.2, as can be seen by straightfoward application of the triangle inequality and the Hölder inequality. This is not the case, if we only assume that $s(\cdot,\omega)$ continuous for each $\omega\in\Omega$.
@@
Let $y\in\mathcal{Y}$ and $\|\cdot\|:\mathcal{E}\to[0,\infty)$ be a continuous norm. Consider its representation $\|\cdot\|_{\tau}$ in a local trivialization around $y\in\mathcal{Y}$. Then for every $\varepsilon>0$ there is a neighborhood $U$ of $y$, such that
$(1-\varepsilon)\|e_{\tau}\|_{\tau,y}\leq\|e_{\tau}\|_{\tau,\hat{y}}\leq(1+\varepsilon)\|e_{\tau}\|_{\tau,y}\quad\forall\hat{y}\in U,e_{\tau}\in\mathbb{E}.$
@@
To define a metric (in the sense of topology) on a manifold $\mathcal{X}$ we will consider a continuous norm on the tangent bundle $T\mathcal{X}$. The following integral over a (piecewise) $C^{1}$-curve $\alpha:[a,b]\rightarrow\mathcal{X}$ is well defined, because the integrand is (piecewise) continuous on $[a,b]$:
@@
The value of this integral can be interpreted as the length of the curve $\alpha$. The following derivations are analogous to the construction of a Riemannian distance (cf. e.g. [14, VII §6]), and generalize them from the case of a Hilbert space norm to the case of a Banach space norm.
@@
In the following we will always assume that $\mathcal{X}$ is connected. Then most of the axioms of a metric follow readily. Symmetry $d(x,y)=d(y,x)$ and non-negativity are obvious from the definition. It is known from topology that on a connected manifold, two points can always be connected by a piecewise smooth curve [13], and thus $d(x,y)<\infty\ \forall x,y\in\mathcal{X}$. The triangle inequality follows straightforwardly from the fact that the length of curves adds up if they are concatenated.
It thus remains to show positive definiteness of $d$, which is the only property of a metric that requires fibrewise uniform continuity. To this end, we first prove some results which relate the metric to the norm, locally. Let $(U,\phi)$ be a local chart at $x_{0}\in\mathcal{X}$ and $\Phi:\pi^{-1}(U)\to U\times\mathbb{X}$ be the corresponding natural vector bundle chart on the tangent bundle $\pi:T\mathcal{X}\to\mathcal{X}$. We define the open ball with radius $r>0$ around 0 with respect to the local norm $\|\cdot\|_{x_{0}}$ (with representation $\|\cdot\|_{\Phi,x_{0}}$) as
$rB_{x_{0}}:=\{v\in\mathbb{X}:\|v\|_{\Phi,x_{0}}<r\}.$
Note that in this special case the model space $\mathbb{E}$ of the vector bundle coincides with the model space $\mathbb{X}$ of the manifold. In particular, applying $\|\cdot\|_{\Phi,x_{0}}$ to chart representations of elements of $\mathcal{X}$ is well defined.
@@
Let $x_{0}\in\mathcal{X}$ and $(U,\phi)$ a local chart, such that $\phi(x_{0})=0$. Let $x,y\in U$ with representatives $x_{\phi}$ and $y_{\phi}$ in the chart. Let $\varepsilon>0$. Then there exists a radius $r>0$ such that for every $C^{1}$-curve $\alpha:[a,b]\to\mathcal{X}$ connecting $x$ and $y$ the following holds:
@@
    If the image $\alpha_{\phi}:[a,b]\to\mathbb{X}$ lies in the open ball $rB_{x_{0}}$, in particular $x_{\phi},\;y_{\phi}\in rB_{x_{0}}$, then the following properties hold:
@@
    If $\alpha_{\phi}$ is leaving the ball $rB_{x_{0}}$, i.e. $\alpha_{\phi}(a)=x_{\phi}\in rB_{x_{0}}$ but $\alpha_{\phi}\not\subset rB_{x_{0}}$, then it holds that
    $L(\alpha)\geq(1-\varepsilon)(r-\|x_{\phi}\|_{\Phi,x_{0}}).$
@@
$(1-\varepsilon)\|v\|_{\Phi,x_{0}}\leq\|v\|_{\Phi,\xi}\leq(1+\varepsilon)\|v\|_{\Phi,x_{0}}\;\forall\xi\in V,v\in\mathbb{X}.$
The image of $V$ in the chart contains a $\|\cdot\|_{\Phi,x_{0}}$-ball $rB_{x_{0}}$ around $\phi(x_{0})=0$ with radius $r=r(\varepsilon)>0$. Let $x,y\in U$ and $\alpha:[a,b]\to\mathcal{X}$ be a $C^{1}$-curve connecting $x$ and $y$.
@@
    Assume that the image $\alpha_{\phi}$ lies in $rB_{x_{0}}$. Using that the curve integral in vector spaces is always greater than or equal to the distance of the end points, we can estimate the length of $\alpha$ by
@@
    Choosing the connecting line of $x_{\phi}$ and $y_{\phi}$ for $\alpha_{\phi}$ this yields $d(x,y)\leq(1+\varepsilon)\|x_{\phi}-y_{\phi}\|_{\Phi,x_{0}}$.
@@
    If $\alpha_{\phi}$ leaves the open ball $rB_{x_{0}}$, there must be a first intersection point $s_{\phi}=\alpha_{\phi}(c)$, $c\in[a,b]$, with the sphere $r\mathbb{S}_{x_{0}}\subset\mathbb{X}$ around 00. Let $\widetilde{\alpha}_{\phi}$ be the part of $\alpha_{\phi}$ connecting $x_{\phi}$ and $s_{\phi}$. Then by the norm equivalence and the inverse triangle inequality:
@@
    $\geq(1-\varepsilon)(\|s_{\phi}\|_{\Phi,x_{0}}-\|x_{\phi}\|_{\Phi,x_{0}})=(1-\varepsilon)(r-\|x_{\phi}\|_{\Phi,x_{0}}).$
    Thus, we get $L(\alpha)\geq(1-\varepsilon)(r-\|x_{\phi}\|_{\Phi,x_{0}})$.
@@
Let $x_{0}\in\mathcal{X}$ and $(U,\phi)$ a local chart and let $\varepsilon>0$. Then there is a neighborhood $W\subseteq U$ of $x_{0}$ such that
$(1-\varepsilon)\,\|x_{\phi}-y_{\phi}\|_{\Phi,x_{0}}\leq d(x,y)\leq(1+\varepsilon)\,\|x_{\phi}-y_{\phi}\|_{\Phi,x_{0}}\ \forall x,y\in W,$
@@
W.l.o.g. assume $\phi(x_{0})=0$. By Lemma 5.5, we directly obtain
$d(x,y)\leq(1+\varepsilon)\|x_{\phi}-y_{\phi}\|_{\Phi,x_{0}}\;\forall x,y\in\phi^{-1}(rB_{x_{0}}),$
@@
$\alpha_{\phi}\subset rB_{x_{0}}$ $\Rightarrow L(\alpha)\geq(1-\varepsilon)\|x_{\phi}-y_{\phi}\|_{\Phi,x_{0}},$
$\alpha_{\phi}\not\subset rB_{x_{0}}$ $\Rightarrow L(\alpha)\geq(1-\varepsilon)(r-\|x_{\phi}\|_{\Phi,x_{0}}).$
@@
$d(x,y)\geq(1-\varepsilon)\cdot\min\{\|x_{\phi}-y_{\phi}\|_{\Phi,x_{0}},r-\|x_{\phi}\|_{\Phi,x_{0}}\}.$ (11)
Choosing $W:=\phi^{-1}(\frac{r}{3}B_{x_{0}})$ yields $d(x,y)\geq(1-\varepsilon)\|x_{\phi}-y_{\phi}\|_{\Phi,x_{0}}$ and thus our claim. ∎
@@
Let $\mathcal{X}$ be a connected $C^{1}$-Banach manifold with the Hausdorff property. Then $d$ defines a metric on $\mathcal{X}$. Moreover, the induced topology coincides with the topology on $\mathcal{X}$ induced by the atlas.
@@
Since $\mathcal{X}$ satisfies the Hausdorff property and $x\neq y$ holds, there is a $\|\cdot\|_{\mathbb{X}}$-ball, and by equivalence of norms also a $\|\cdot\|_{\Phi,x}$-ball $rB_{x}$ around 00 with radius $r>0$ such that $y\not\in\phi^{-1}(rB_{x})$. Possibly decreasing $r$ first, we obtain by (11)
@@
Moreover, for $y\in\phi^{-1}(r\mathbb{S}_{x})$ we obtain by Prop. 5.6
@@
Now consider a retraction $R:T\mathcal{X}\rightarrow\mathcal{X}$. We will relate norms of its preimages on the tangent spaces, to the distance of its images on $\mathcal{X}$, measured on the corresponding metric. First of all, we establish existence and smoothness of an inverse: